OpenStudy (anonymous):

Using elimination to solve the system 2x + y = 8 and x + y = 4, which of the following could be the first step? A. 2x + y = 8--------> 2x + y = 8 (x + y = 4)(-1)---> -x - y = -4 B. 2x + (4 - x) = 8 C. 2x + y = x + y D. 3x + 2y = 12 E. x = 4

OpenStudy (anonymous):

I said it was A...

OpenStudy (anonymous):

B

OpenStudy (anonymous):

A-B

OpenStudy (anonymous):

What does that mean?

OpenStudy (anonymous):

basically you take away the 2nd equation from the first equation

OpenStudy (anonymous):

to eliminate y

OpenStudy (anonymous):

then find x

OpenStudy (anonymous):

So its E?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

what level question is this?

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

what year? gcse year 9, 10?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

oh ok thats good

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

i just wanted to know... I remember doing that question for a gcse paper long time ago

OpenStudy (anonymous):

2x+y=8 (x+y=4) *-1 2x+y=8 -x-y=-4 The FIRST step would be A

OpenStudy (anonymous):

Wait.. E or A?

OpenStudy (anonymous):

You can't get to x=4 using elimination without first doing step A or an equivalent step

OpenStudy (anonymous):

You could also multiply the 1st equation by -1, but that's not an option